# HSSBench: Benchmarking Humanities and Social Sciences Ability for Multimodal Large Language Models

### v1

@@ 1 Introduction
Multimodal large language models (MLLMs) achiam2023gpt ; team2024gemini have demonstrated remarkable performance across a wide range of tasks, recently achieving or even exceeding human-level capabilities in many of them. As the performance of MLLMs continues to improve, conducting a comprehensive evaluation of their capabilities has become increasingly essential. In recent times, numerous benchmarks for evaluating MLLMs have emerged hendrycks2020measuring ; yue2024mmmu ; liu2024mmbench ; saikh2022scienceqa ; zhang2024cmmmu . These tasks are designed to assess the models’ ability to jointly understand and reason across multiple modalities, such as images and text, from various perspectives.
Specifically, most multimodal benchmarks are designed either from a general perspective liu2024mmbench ; liu2023visualinstructiontuning ; meng2024vgavisionguiassistant ; han2024instinctivebiasspuriousimages ; qian2024easyfoolmultimodalllms or with a focus on scientific disciplines such as mathematics wang2024measuringmultimodalmathematicalreasoning ; lu2024mathvistaevaluatingmathematicalreasoning science li2023scigraphqalargescalesyntheticmultiturn ; liang2024scemqascientificcollegeentrance , and programming song2025csbench . Unlike STEM fields that employ "vertical reasoning": a focused sequential process using logical deduction and experimental analysis to arrive at singular correct answers, the Humanities and Social Sciences (HSS) emphasize "horizontal reasoning," requiring connections across different contexts and generating multiple valid interpretations rather than single solutions. This fundamental difference stems from the inherent attributes of these disciplines: While STEM fields utilize relatively fixed symbolic systems with standardized reading sequences, HSS disciplines feature symbol systems deeply rooted in regional cultures with meanings that require historical-cultural context interpretation. Furthermore, STEM knowledge can be iteratively developed through logical deduction and experimental analysis, whereas HSS knowledge verification relies on more complex pathways involving cross-referencing literature and expert consensus, with strong dependencies on real-world information.
@@ 2 HSSBench
The worldview within the HSS is expansive and lacks unified definitions. Our dataset focuses on six high-interest categories within the HSS domain: geography, art, culture, social sciences, history, and economics. Textual data are the primary medium for disseminating knowledge in the HSS domain, with almost no data in pictorial form. This presents significant challenges for experts in the construction of datasets. For this purpose, we designed and developed a VQA Generation Pipeline (VGP) to generate the dataset. After engaging experts for annotation, we designed a data construction agent based on their annotation logic, enabling us to generate a sufficient volume of data.
@@
2.1 Dataset Preparation, Stage \@slowromancapi@
@@
Figure 3: Our pipeline to build HSSBench.
Data Selection and Collection with Experts
@@
Networked Information-Aggregation Agent
@@
2.2 Dataset Construction, Stage \@slowromancapii@
@@
We require each question to be presented in a multiple-choice format, with one correct answer and three distractors that were plausible but ultimately incorrect. At this stage, experts have two main responsibilities: revising existing questions and creating entirely new ones. 1) When revising existing questions, experts optimize multiple-choice items, encompassing both the original ones provided in the materials and those requiring revision after failing validation tests. Experts carefully review the relevant textual and visual content, adjusting the question stems and answer choices by integrating real-world knowledge. The goal is to strengthen the connection between the questions and the accompanying images while improving the plausibility of the incorrect options. 2) For creating new questions, experts generate items based on the given high-quality texts and images. These new questions are expected to demand identification of knowledge points from the images and complex reasoning based on real-world understanding to arrive at the correct answers. Following these principles, experts produce a set of high-quality questions that fully utilize the provided multimodal materials to meet the demands of model performance evaluation.
@@ 2.3 Validation, Stage \@slowromancapiii@
This stage aims to eliminate duplicate questions and verify the strength of the correlation between the visual content and the corresponding text. First, the Validation Agent calculates the textual similarity between questions and filters out highly redundant ones based on predefined criteria to ensure the diversity of the dataset. Next, for all constructed data, it is necessary to ensure that 1) without providing an image, the question cannot be correctly answered based solely on the text, and 2) without providing a question, the image alone cannot lead to the correct answer. This requirement stems from our aim to scrutinize the model’s capabilities from a multimodal perspective. If a single modality, text or images, suffices to answer a question, it compromises the quality of the dataset and its ability to comprehensively evaluate the performance of MLLMs. To meet these two requirements, the Validation Agent assesses the extent to which the question depends on the image. If the image is deemed unnecessary, the question is sent back to the Stage \@slowromancapii@ for revision; if it still fails to meet the requirements after multiple iterations, the question is discarded.
@@ 2.4 Implementation Details
In our VGP, we leverage GPT-4o and GPT-4.1 due to their outstanding capabilities. Throughout all stages of LLM, we employed the Chain-of-Thought prompting strategy. For more detailed information on VGP, please refer to Appendix A.
@@ 2.5 Data Statistics
Our dataset comprises a total of 13,152 questions spanning six categories and 45 types. It supports evaluation in six different languages. The statistical data of HSSBench is illustrated in Figure 4, with a detailed description of the dataset provided in Appendix B.
@@ 3.1 Experimental Setup
For open-source models, we selected Qwen2.5-VL-3B/7B/32B/72B Qwen2.5-VL , Qwen2-VL-72B, QVQ-72B-Preview Qwen2-VL , Deepseek-VL2-Tiny wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels , MiniCPM-o-2.6 yao2024minicpm , mPLUG-Owl3-2B/7B ye2024mplugowl3longimagesequenceunderstanding , Llava-onevision-7B li2024llavaonevisioneasyvisualtask , Llama3-llava-next-8b li2024llavanext-strong , InternVL3-8B zhu2025internvl3exploringadvancedtraining , InternVL2.5-8B-MPO chen2024expanding , Phi-3.5-Vision-Instruct abdin2024phi3technicalreporthighly , Janus-Pro chen2025janusprounifiedmultimodalunderstanding , Llava1.5 liu2024improvedbaselinesvisualinstruction . For closed-source commercial models, we utilized GPT-4o/4.1/4.1-mini/4.1-nano openai2024gpt4ocard .
@@
1) We conducted evaluations using two types of prompts to examine the MLLMs performance under different prompting strategies. One approach involved prompting the model to directly output the answer without any intermediate reasoning, while the other required the model to generate a chain of thought before providing the final answer. 2) Additionally, we tested each question in six different language versions to investigate how the MLLMs’ performance varies across languages for the same question. 3) Furthermore, we optimized the question formats by designing two types of prompts for each item: multiple-choice and open-ended questions. This allowed us to assess whether the model can correctly answer questions without any explicit hints. 4) For the HSSBench evaluation, we employed two assessment methods and selected GPT-4o as a representative closed-source model and Qwen2.5-7B as a representative open-source model.A detailed comparative analysis of both models is provided in Appendix C.5.
@@ 3.2 Results
Table 1 presents the overall results of various open-source and closed-source MLLMs evaluated on HSSBench in the context of English language tests. Ct. and Dt. represent two types of prompts: Chain-of-Thought (COT) prompts and direct response prompts. C. and O. denote question types: multiple-choice and open-ended questions. "Human" refers to the average final scores achieved by experts in various fields. Detailed experimental results for Dt. and are presented in Appendix C.1.
@@
The experimental results highlight the performance differences between the various models. Of the open-source models, Qwen2.5-VL-72B-Instruct delivers the highest performance, although it still falls short of surpassing closed-source models in open-ended question tests. The GPT-4.1 series achieves state-of-the-art performance in most tasks, with a particularly impressive accuracy of 39.97% in open questions, almost double that of other closed-source models. In contrast, some open-source models show considerably lower performance.
Although data in the HSS domain was generally considered common knowledge, our study demonstrates that from a multimodal perspective, these models require more diverse training data and targeted fine-tuning to effectively address related tasks. We also observe that model performance varies by language. Multilingual evaluation results are provided in the Appendix C.3.
@@
Open-source LLM (Scale <<< 10B)
@@
Open-source LLM (Scale >>> 10B)
@@
Table 1: Scores (%) of MLLMs on HSSBench (EN-\@slowromancapi@). The highest and second highest scores are marked in blue and green, respectively.
@@
Among the six categories evaluated on HSSBench (EN-\@slowromancapi@), the economic-related tasks consistently emerge as the most challenging for the models. The average score of all models in this category is the lowest, indicating that addressing economic problems requires a deep understanding of various economic theories and the ability to apply them in complex reasoning. However, current open-source MLLMs still exhibit significant deficiencies in these aspects. In contrast, closed-source models perform exceptionally well in economically related tasks, and this advantage may stem from their exposure to large amounts of high-quality training data in the economic domain.
@@ 3.3 Qualitative Analysis
Table 1 demonstrates that the performance of various MLLMs on HSSBench using direct answer prompts and COT prompts exhibits subtle differences. COT prompts do not always yield positive effects, and certain models perform better with direct answer prompts. This suggests that extending the reasoning process might mislead the models.
In HSS tasks, the COT method exhibits limitations for certain models. It not only fails to deepen the models’ understanding but may also cause them to focus on irrelevant information, thereby interfering with attention to critical knowledge. Specifically, COT prompts exacerbate hallucination issues in certain models, where reasoning flaws in textual analysis and misinterpretation of visual inputs lead to the generation of incorrect background knowledge during step-by-step analysis. This causes reasoning to deviate from the correct answer. For example, when tackling geographic questions, models often struggle to accurately interpret spatial elements such as location markers and contour lines—visual features primarily composed of points and lines—resulting in analytical outcomes that contradict the actual image content.
@@
We selected questions from the dataset that can be reformulated as open-ended to evaluate model performance without the guidance of provided answer options. The results indicate that HSS tasks pose significant challenges for current MLLMs, with only a few models achieving accuracy above 15% in the open-ended setting. This result aligns with our expectations: even experts find it difficult to provide the correct answer without predefined options and background knowledge.
The presence of answer options offers models valuable prior knowledge, effectively constraining the scope of possible answers. In contrast, open-ended questions remove this constraint, causing the model’s reasoning to become highly divergent and often leading to answers that deviate substantially from the ground truth. This suggests that models primarily rely on shallow visual attributes, such as object size, type, or motion, while overlooking deeper symbolic information embedded in cultural context, academic theories, or spatiotemporal cues.
These observations highlight two key limitations of current MLLMs: first, a lack of comprehensive understanding and memory of real-world knowledge; second, difficulty in effectively integrating visual features with textual information. Moreover, the performance degradation observed when introducing chain-of-thought prompting further underscores the models’ struggles with internalizing and retrieving complex knowledge.
@@
In multiple-choice settings, the presence of answer options inevitably provides additional cues that can influence model performance. To further investigate whether models truly understand the questions rather than relying on superficial cues, we introduced confusing options, specifically, the choice "None of the above," while keeping the correct answer unchanged. This experiment was conducted on a subset of 150 samples randomly selected from each category. The results are illustrated in Table 5.
For lower-performing models, the introduction of confusing options led to a noticeable decline in performance. Upon closer examination of the models’ reasoning processes after adding these distractors, two main patterns emerged. In some cases, models were directly swayed by the confusing option and selected "None of the above" incorrectly. More intriguingly, in other cases, although the models did not explicitly choose the confusing option, their reasoning became less coherent.
This behavior indicates that current models struggle to accurately assess the relative credibility of different answer choices and to make reliable judgments when faced with conflicting information. Consequently, their performance deteriorates in the presence of carefully designed distractors.
@@
To investigate the impact of visual information loss on the performance of MLLMs, we designed two complementary experiments. In the first experiment, we used GPT-4.1 to generate detailed textual descriptions of the images. In the second experiment, we invited domain experts to produce comprehensive and precise annotations for each image. The experimental results are summarized in Table 6.
Specifically, we converted the content of images into textual descriptions and provided only these texts, along with the questions and answer options, to the models. The accuracy of most models declined compared to when the full images were available, indicating that direct visual input contains critical information that is partially lost in the text conversion process.
However, considering that even when models receive the full images, they may not fully comprehend certain fine-grained visual details, we further involved domain experts to produce detailed annotations of the images. These expert annotations distilled key visual information into precise textual descriptions. When provided with these expert-annotated texts instead of the images, the models’ accuracy notably improved relative to the plain text conversion scenario, despite the absence of the original images.
Interestingly, some models even surpassed their previous upper-bound performance in domains such as Culture and Social Sciences. This improvement suggests that expert annotations effectively guide the models to focus on critical visual cues by translating them into semantically rich and relevant textual information, thereby facilitating stronger connections with domain knowledge. These findings corroborate our earlier observations on open-ended questions, highlighting inherent limitations in current MLLMs’ abilities to retrieve and interpret visual information comprehensively.
4 Limitations
Our HSSBench is designed to evaluate MLLMs’ capabilities through HSS reasoning tasks. The type within our dataset reflects our effort to collect data from as many perspectives as possible. However, we acknowledge that the scope of knowledge covered is inherently limited by the contributors’ expertise and cannot encompass all aspects of HSS. Moreover, HSS currently lacks a universally accepted definition. Our evaluation is based on six major categories that, while comprehensive, do not cover the full spectrum of HSS-related fields.
In addition, some data points involve recent social events and the latest economic knowledge. We recognize that such data are time sensitive. Although the answers are accurate at present, future developments may lead to different interpretations. For these items, we impose temporal and spatial constraints to ensure their correctness within the current context.
Although these limitations are relevant, we believe that they extend beyond the scope and original intent of our dataset. We hope that future work will address these challenges in more detail.

### v2

@@ 1 Introduction
Multimodal large language models (MLLMs) Achiam et al. (2023); Team et al. (2024) have demonstrated remarkable performance across a wide range of tasks, recently achieving or even exceeding human-level capabilities in many of them. As the performance of MLLMs continues to improve, conducting a comprehensive evaluation of their capabilities has become increasingly essential. In recent times, numerous benchmarks for evaluating MLLMs have emerged Hendrycks et al. (2020); Yue et al. (2024); Liu et al. (2024b); Saikh et al. (2022); Zhang et al. (2024). These tasks are designed to assess the models’ ability to jointly understand and reason across multiple modalities, such as images and text, from various perspectives.
Specifically, most multimodal benchmarks are designed either from a general perspective Liu et al. (2024b; 2023); Meng et al. (2024); Han et al. (2024); Qian et al. (2024) or with a focus on scientific disciplines such as mathematics Wang et al. (2024a); Lu et al. (2024) science Li and Tajbakhsh (2023); Liang et al. (2024), and programming Song et al. (2025). Unlike STEM fields that employ "vertical reasoning": a focused sequential process using logical deduction and experimental analysis to arrive at singular correct answers, the Humanities and Social Sciences (HSS) emphasize "horizontal reasoning," requiring connections across different contexts and generating multiple valid interpretations rather than single solutions. This fundamental difference stems from the inherent attributes of these disciplines: While STEM fields utilize relatively fixed symbolic systems with standardized reading sequences, HSS disciplines feature symbol systems deeply rooted in regional cultures with meanings that require historical-cultural context interpretation. Furthermore, STEM knowledge can be iteratively developed through logical deduction and experimental analysis, whereas HSS knowledge verification relies on more complex pathways involving cross-referencing literature and expert consensus, with strong dependencies on real-world information.
@@ 2 HSSBench
The worldview within the HSS is expansive and lacks unified definitions. Our dataset focuses on six high-interest categories within the HSS domain: geography, art, culture, social sciences, history, and economics. Textual data remain the primary medium for disseminating knowledge in the HSS domain, while pictorial data, though less abundant and more challenging to collect, provide valuable complementary information. This presents significant challenges for experts in the construction of datasets. For this purpose, we designed and developed a VQA Generation Pipeline (VGP) to generate the dataset. After engaging experts for annotation, we designed a data construction agent based on their annotation logic, enabling us to generate a sufficient volume of data.
@@
2.1 Dataset Preparation, Stage I
@@
Data Selection and Collection with Experts.
@@
Networked Information-Aggregation Agent.
@@
2.2 Dataset Construction, Stage II
@@
At this stage, experts have two main responsibilities: revising existing questions and creating entirely new ones. 1) When revising existing questions, experts optimize multiple-choice items, encompassing both the original ones provided in the materials and those requiring revision after failing validation tests. Experts carefully review the relevant textual and visual content, adjusting the question stems and answer choices by integrating real-world knowledge. The goal is to strengthen the connection between the questions and the accompanying images while improving the plausibility of the incorrect options. 2) For creating new questions, experts generate items based on the given high-quality texts and images. These new questions are expected to demand identification of knowledge points from the images and complex reasoning based on real-world understanding to arrive at the correct answers. Following these principles, experts produce a set of high-quality questions that fully utilize the provided multimodal materials to meet the demands of model performance evaluation.
@@ 2.3 Validation, Stage III
This stage aims to eliminate duplicate questions and verify the strength of the correlation between the visual content and the corresponding text. First, the Validation Agent calculates the textual similarity between questions and filters out highly redundant ones based on predefined criteria to ensure the diversity of the dataset. Next, for all constructed data, it is necessary to ensure that 1) without providing an image, the question cannot be correctly answered based solely on the text, and 2) without providing a question, the image alone cannot lead to the correct answer. This requirement stems from our aim to scrutinize the model’s capabilities from a multimodal perspective. If a single modality, text or images, suffices to answer a question, it compromises the quality of the dataset and its ability to comprehensively evaluate the performance of MLLMs. To meet these two requirements, the Validation Agent assesses the extent to which the question depends on the image. If the image is deemed unnecessary, the question is sent back to the Stage II for revision; if it still fails to meet the requirements after multiple iterations, the question is discarded.
@@ 2.4 Implementation Details
In our VGP, we leverage GPT-4o and GPT-4.1 due to their outstanding capabilities. Throughout all stages of LLM, we employed the Chain-of-Thought(COT) prompting strategy. For more detailed information on VGP, please refer to Appendix A.
@@ 2.5 Data Statistics
HSSBench consists of 13,152 multiple-choice questions distributed across six major categories—Economy, Art, Culture, Social Sciences, History, and Geography—and further divided into 45 specialized subtypes.
We require each question to be presented in a multiple-choice format with one correct answer and several distractors that are plausible but ultimately incorrect. Although some original questions in the HSS domain naturally have multiple correct answers, these multi-answer questions have been reformulated into single-answer questions for consistency. This was achieved by combining the multiple correct options into a single choice. This approach preserves the original content and complexity of the questions while conforming to a uniform single-answer multiple-choice format.
It supports evaluation in six different languages. Our initial data came from multiple countries and languages, with questions originally created by domain experts in their source languages (examples in Appendix D). We then used LLM-based translation models to translate the questions into five other languages. All translations were carefully reviewed and validated by bilingual experts to ensure linguistic accuracy and cultural appropriateness, maintaining semantic consistency while respecting cultural nuances.
The statistical data of HSSBench is illustrated in Figure 4, with a detailed description of the dataset provided in Appendix B. Appendix B.1 reports human expert accuracy across all categories. Appendix B.2 provides an in-depth taxonomy of categories and subtypes. Appendix B.3 also includes an analysis of the most frequent content words per category. Appendix B.4 details the composition of the dataset in terms of contributions of human experts versus automated agents, together with quality validation through model accuracy comparisons.
@@ 3.1 Experimental Setup
For open-source models, we selected Qwen2.5-VL-3B/7B/32B/72B Bai et al. (2025), Qwen2-VL-72B Wang et al. (2024c), QVQ-72B-Preview Wang et al. (2024b), Deepseek-VL2-Tiny Wu et al. (2024c), MiniCPM-o-2.6 Yao et al. (2024), mPLUG-Owl3-2B/7B Ye et al. (2024), Llava-onevision-7B Li et al. (2024b), Llama3-llava-next-8b Li et al. (2024a), InternVL3-8B Zhu et al. (2025), InternVL2.5-8B-MPO Chen et al. (2024b), Phi-3.5-Vision-Instruct Abdin et al. (2024), Janus-Pro Chen et al. (2025), Llava1.5 Liu et al. (2024a). For closed-source commercial models, we utilized GPT-4o/4.1/4.1-mini/4.1-nano OpenAI et al. (2024).
@@
1) We conducted evaluations using two types of prompts to examine the MLLMs performance under different prompting strategies. One approach involved prompting the model to directly output the answer without any intermediate reasoning, while the other required the model to generate a COT before providing the final answer. 2) Additionally, we tested each question in six different language versions to investigate how the MLLMs’ performance varies across languages for the same question. 3) Furthermore, we optimized the question formats by designing two types of prompts for each item: multiple-choice and open-ended questions. This allowed us to assess whether the model can correctly answer questions without any explicit hints. Detailed prompt settings are provided in Appendix C.8. 4) For the HSSBench evaluation, we employed two assessment methods and selected GPT-4o as a representative closed-source model and Qwen2.5-7B as a representative open-source model. A detailed comparative analysis of both models is provided in Appendix C.9.
@@ 3.2 Results
Table 1 presents the overall results of various open-source and closed-source MLLMs evaluated on HSSBench in the context of English language tests. Ct. and Dt. represent two types of prompts: COT prompts and direct response prompts. C. and O. denote question types: multiple-choice and open-ended questions. "Human" refers to the average final scores achieved by experts in various fields. Detailed experimental results for Dt. and are presented in Appendix C.1.
@@
The experimental results highlight the performance differences between the various models. Of the open-source models, Qwen2.5-VL-72B-Instruct delivers the highest performance, although it still falls short of surpassing closed-source models in open-ended question tests. The GPT-4.1 series achieves state-of-the-art performance in most tasks, with a particularly impressive accuracy of 39.97% in open questions, almost double that of other closed-source models. In contrast, some open-source models show considerably lower performance. We also observe that model performance varies by language. Multilingual evaluation results are provided in the Appendix C.2.
@@
Open-source LLM (Scale << 10B)
@@
Open-source LLM (Scale >> 10B)
@@
Table 1: Scores (%) of MLLMs on HSSBench (EN-I). The highest and second highest scores are marked in blue and green, respectively.
@@
Among the six categories evaluated on HSSBench (EN-I), the economic-related tasks consistently emerge as the most challenging for the models. The average score of all models in this category is the lowest, indicating that addressing economic problems requires a deep understanding of various economic theories and the ability to apply them in complex reasoning. However, current open-source MLLMs still exhibit significant deficiencies in these aspects. In contrast, closed-source models perform exceptionally well in economically related tasks, and this advantage may stem from their exposure to large amounts of high-quality training data in the economic domain.
@@ 3.3 Qualitative Analysis
Table 1 shows subtle differences in MLLMs’ performance on HSSBench when using direct answer versus COT prompting. Notably, COT does not always help; some models perform better with direct answers, indicating that longer reasoning can mislead them.
Specifically, COT prompts exacerbate hallucination issues in certain models, where reasoning flaws in textual analysis and misinterpretation of visual inputs lead to the generation of incorrect background knowledge during step-by-step analysis. This causes reasoning to deviate from the correct answer. For example, when tackling geographic questions, models often struggle to accurately interpret spatial elements such as location markers and contour lines—visual features primarily composed of points and lines—resulting in analytical outcomes that contradict the actual image content.
@@
We reformulated some questions as open-ended to test models without answer options. Results reveal that HSS tasks remain very challenging: only a few models exceed 15% accuracy. This matches expectations, as even experts find these questions difficult without options or background cues.
Answer choices provide prior knowledge that narrows the answer space. Without them, models’ reasoning becomes highly divergent, often drifting far from correct answers. This suggests models mainly rely on shallow visual features (e.g., size, type, motion) but miss deeper symbolic information like cultural context or spatiotemporal cues.
These findings highlight two key limitations: insufficient real-world knowledge and weak integration of visual and textual modalities. The performance drop with COT further reflects models’ struggles to internalize and retrieve complex knowledge. Improving model capabilities on HSS tasks remains a critical challenge.
@@
To test if models truly understand questions rather than exploiting options, we added confusing choices like “None of the above” while keeping the correct answer unchanged, sampling 150 questions per category (results in Table 11, Appendix C.3).
Lower-performing models’ accuracy dropped noticeably. Analysis revealed two patterns: some models were misled to select the confusing option; others, though not choosing it, showed incoherent reasoning, failing to filter out wrong choices and sometimes excluding the correct answer altogether. This indicates models struggle to assess the relative credibility of options and make reliable judgments, exposing their vulnerability to distractors and raising concerns about robustness.
@@
To investigate the impact of visual information loss on the performance of MLLMs, we designed two complementary experiments. In the first experiment, we used GPT-4.1 to generate detailed textual descriptions of the images. In the second experiment, we invited domain experts to produce comprehensive and precise annotations for each image (results in Table 12, Appendix C.4).
Most models’ accuracy dropped when images were replaced by GPT-generated texts, showing that direct visual input contains critical details lost in conversion. However, with expert annotations, accuracy improved noticeably despite no image access. Some models even surpassed previous performance ceilings in Culture and Social Sciences, as expert texts helped focus on crucial visual cues and better link to domain knowledge.
These results confirm earlier observations that current MLLMs have inherent limitations in retrieving and understanding visual information fully. They underscore the importance of improving models’ abilities to extract and integrate key visual features for HSS tasks.
Comparative Analysis of HSSBench with Related Benchmarks.
We compared HSSBench with other benchmarks that include some HSS data, such as CMMMU, MME, and MMMU, as well as with STEM benchmarks (details in Appendix C.5 and C.6).
Across HSS benchmarks, relative model performance trends are consistent, but HSSBench yields lower accuracy, indicating it is more challenging and better captures HSS complexity. Compared to STEM benchmarks, models perform much better on STEM tasks, highlighting the unique difficulties of HSS domains that require nuanced cultural and social understanding. These comparisons demonstrate HSSBench’s value in driving progress on underexplored HSS challenges.
Evaluation of Retrieval-Augmented Generation on HSSBench.
We tested Retrieval-Augmented Generation (RAG) by integrating external knowledge from Wikipedia and HSS documents with several smaller MLLMs under direct and COT prompting (results in Tables 15 and 16, Appendix C.7).
Contrary to expectations, RAG did not consistently improve accuracy. Some modest gains appeared in specific domains or models, but overall performance was often similar or worse than direct prompting alone. This suggests that general retrieval corpora and simple integration methods are insufficient for the nuanced knowledge HSS tasks demand.
The limited effectiveness of RAG in supporting MLLMs for complex multi-hop HSS tasks stems from misalignment between general-purpose corpora and HSS knowledge, as well as deficiencies in current retrieval and integration methods. Moreover, MLLMs struggle to internalize and transfer newly retrieved domain knowledge. These challenges highlight the need for specialized HSS retrieval resources, RAG techniques tailored to HSS, and enhanced mechanisms for MLLMs to acquire, transfer, and apply HSS-specific knowledge.